A triangular lamina in the xy -plane such that its vertices are (0,0), (0,1) and (1,0). Suppose that the density function of the lamina is defined by p(x,y)=120xy. What is the total mass of the lamina and the center of gravity?
1
Expert's answer
2020-05-25T21:13:27-0400

We assume all units are in S.I unit.

Let's draw the rough sketch of the given triangular lamina.




Given, density of the triangular lamina

"\rho(x,y)=120xy"

Since, by definition of density,

"\frac{dm}{dA}=\rho(x,y)\\
\implies m=\int_{0}^{1}\int_{0}^{1}\rho(x,y)dA=\int_{0}^{1}\int_{0}^{1}\rho(x,y)dxdy\\
\implies m=\int_{0}^{1}\int_{0}^{1}120xy \:dxdy\\\implies m=120\int_{0}^{1}\bigg (y\int_{0}^{1}xdx\bigg)dy=30Kg"

Let's denote the center of gravity by "(X_G,Y_G)" .

Now,

"X_G=\frac{1}{m}\int_{0}^{1}\int_{0}^{1}x\rho(x,y)dxdy\\
\implies X_G=\frac{1}{m}\int_{0}^{1}\int_{0}^{1}x(120xy)dxdy\\
\implies X_G=\frac{120}{m}\int_{0}^{1}\int_{0}^{1}x^2ydxdy\\
\implies X_G=\frac{120}{m}\int_{0}^{1}x^2\bigg(\int_{0}^{1}ydy\bigg)dx\\
\implies X_G=\frac{60}{m}\int_{0}^{1}x^2dx\\
\implies X_G=\frac{20}{m}=\frac{2}{3}m"

Similarly,

"Y_G=\frac{1}{m}\int_{0}^{1}\int_{0}^{1}y\rho(x,y)dxdy"

Clearly, if we replace "x\leftrightarrow y" and vice versa we get exactly similar expression as of "X_G" ,thus we conclude that

"Y_G=\frac{2}{3}m"

Therefore center of gravity of the given triangular lamina is

"(X_G,Y_G)=\bigg(\frac{2}{3},\frac{2}{3}\bigg)"